OpenStudy (anonymous):

Identify the solution set of |x + 3| + 7 equal to 10 and describe the graph in a complete sentence.

OpenStudy (anonymous):

TO get the solution set start with |x + 3| + 7 = 10 Then |x + 3| = 3, so now you have x + 3 = 3 and -(x + 3) = 3 x = 0 and x + 3 = -3 x = 0 and x = -6

OpenStudy (anonymous):

In set notation x is an element of {-6,3}

OpenStudy (anonymous):

The graph will be in the shape of a V with roots at -6 and 3 and a vertex at x=-3/2